By drawing a relevant triangle,prove that

 

tan θ = sin θ / cos θ

and

cos^2 θ + sin^2 θ = 1


1
Expert's answer
2021-06-24T11:43:39-0400

Consider the right-angled triangle ABC, with AB=1, C = 90° and A= θ°.

"\sin \theta =\frac{BC}{AB}=\frac{BC}{1}=BC"


"\cos \theta =\frac{AC}{AB}=\frac{AC}{1}=AC"


"\tan \theta =\frac{BC}{AC}=\frac{\sin \theta}{\cos \theta}"


Pythagorean theorem: "1=AC^2+BC^2=\cos ^2\theta +\sin^2\theta"